(100-100)/(100-100)
= (10square-10square)/(10 (10-10))
= (10+10)(10-10)/10 (10-10)
= (20)/10
= 2
Is it correct or not sir

Found 2 solutions by Fombitz, Theo:
Answer by Fombitz(32388)

No.
You begin with a division by zero (which is undefined) which then leads you to an incorrect solution.

Answer by Theo(13342)

(100 - 100) / (100 - 100) = 0 / 0 = undefined.

following your logic:

1. (100-100)/(100-100)
2. = (10square-10square)/(10 (10-10))
3. = (10+10)(10-10)/10 (10-10)
4. = (20)/10
5. = 2
Is it correct or not sir

number 1 is correct.
number 2 is correct.
number 3 is correct.
number 4 is not correct. *****
the problem with number 4 is as follows:

start with (10+10)(10-10)/10 (10-10) from number 3.

10 + 20 = 20 - this part is true.
10 - 10 = 0.
10 * (10 - 10) = 10 * 0 which is also equal to 0.
you get:
20 * 0 / 0 becomes 0 / 0.

another way to do it is to use the distributive law of multiplication.

start with (10+10)(10-10)/10 (10-10)

this becomes (10 * 10) - (10 * 10) + (10 * 10) + (10 * -10) / ((10 * 10) - (10 * 10))

simplify this to get (100 - 100) + (100 + (-100)) / (100 - 100)

simplify this further to get (100 - 100) + (100 - 100) / (100 - 100)

simplify this further to get (0 + 0) / 0.

simplify this further to get 0 / 0.

all of these machinations are unnecessary, because:

(100 - 100) / (100 - 100) = 0 / 0 = undefined.

you perform the operations indicated within the parentheses first.

100 - 100 = 0

that leads to the rest, because you then get 0 / 0.

if you solve it using the distributive law of multiplication, you get the same answer.